Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
verkindt, grimaud, rolland - 16:52 Wednesday 10 April 2024 (63919) Print this report
Hrec with new actuators models

Today around 13h50 UTC, I have restarted Hrec after changing its configuration to:
- use the new actuators models that were set on the basis of PCal recent updates and calibration measurements done since August 2023
   (those new actuators models were used since April 4 in the test process HrecN)
- use an updated delay of 67 us (instead of 58 us) in the B1 sensing model
- use the bias file /virgoData/Hrec/Hrec_bias_202404.txt that contains the bias and uncertainty on h(t) estimated using the MirNE checkhrec injections done between April 7 and April 10.
  This bias is used to unbias online the h(t), these uncertainties are conservative uncertainties provided with online h(t) as Hrec*U00* channels.
   Plot1 shows in black the bias and in pink the uncertainties versus frequency provided in this file.

On the basis of what HrecN showed during the last days, the use of these updated actuators models should decrease the BNS range by about 1 Mpc (before unbiasing)
and introduce more bias (plot2).
While, the estimation of DCP frequency and optical gain is a bit more accurate (plot3).

Images attached to this report
Comments to this report:
rolland, verkindt - 18:41 Wednesday 10 April 2024 (63925) Print this report

TFMoni has been restarted earlier today (around 14h17 UTC) with updated calibration parameters.

But I found that the DMS flags CALinj was gray.  It is due to some the coherence channels of HREC/HINJ, generated by TFMoni, being corrupted. Restarting TFMoni did not solve the problem. A restart did not solve the problem.

The coherence is computed between Hrec, being 0 since ITF is not locked, and injection channels. We have to check if there is a bug in TFMoni that generate corrupted coherence data when it is  started  while h(t) = 0.

Meanwhile, I have commented the 4 corresponding corrupted checks in CaliMoni, so that the DMS flag is no more impacted for tonight.   We will check the output of TFMoni when the ITF will be relocked and h(t) non null.

Images attached to this comment
Search Help